# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

### v1

@@ Abstract.
Decision-making with information displays is a key focus of research in areas like explainable AI, human-AI teaming, and data visualization. However, what constitutes a decision problem, and what is required for an experiment to be capable of concluding that human decisions are flawed in some way, remain open to speculation. We present a widely applicable definition of a decision problem synthesized from statistical decision theory and information economics. We argue that to attribute loss in human performance to forms of bias, an experiment must provide participants with the information that a rational agent would need to identify the normative decision. We evaluate the extent to which recent evaluations of decision-making from the literature on AI-assisted decisions achieve this criteria. We find that only 6 (17%) of 35 studies that claim to identify biased behavior present participants with sufficient information to characterize their behavior as deviating from good decision-making. We motivate the value of studying well-defined decision problems by describing a characterization of performance losses they allow us to conceive. In contrast, the ambiguities of a poorly communicated decision problem preclude normative interpretation. We conclude with recommendations for practice.
@@ 1. Introduction
Research areas that rely heavily on decision-making would be more coherent if there was a consensus on what a decision is, and how decisions are a distinct form of human behavior compared to tasks that are not considered decisions. Human decisions from prediction displays are often evaluated in studies with the goal of identifying forms of bias: ways in which study participants’ behavior deviates from expectations of good decision-making. Reviewing the range of tasks that get filed under decision-making, it is difficult to ascertain what common task structure makes these studies about decision-making, much less what properties a study needs to have in order for it to be capable of distinguishing human decisions as flawed. Colloquial definitions describe a decision task as a choice between alternatives, or a choice where the stakes are high (such as deciding which of multiple hurricane forecasts is more accurate (Padilla et al., 2015), or whether to screen child welfare cases (Zytek et al., 2021)), but neither isolates a clear set of assumptions that must hold for a judgment to qualify as a decision. Other accounts cast decision-making as a higher level task akin to sensemaking (Dimara and Stasko, 2021). For some, the point of studying human decisions is to compare them to a definition of idealized behavior. However, sometimes researchers define this standard in a way that is common across studies (e.g., a better forecast comes closer to the true probability of the event (Green and Chen, 2019; Fogliato et al., 2021)) while other times different definitions of the same concept are used (e.g., appropriate reliance as agreeing with an AI when it is correct and disagreeing when it’s not (Bansal et al., 2021; Buçinca et al., 2021) versus providing a higher subjective rating of trust when the model is correct (Bussone et al., 2015) versus choosing the better of the human versus the AI prediction to bet on for a future task (Dietvorst et al., 2015)). Sometimes the point is to study seemingly “subjective decisions,” like how people react to music recommendations (Kulesza et al., 2012), or decide whether a house seems to match its valuation (Prabhudesai et al., 2023), where it can seem unclear whether normative interpretations of task performance can be applied. Hence, not only is a more coherent definition of decision needed, as recent survey papers on decision making in visualization and human-centered AI call for (Dimara and Stasko, 2021; Lai et al., 2021), but such a definition must clearly define requirements for identifying forms of bias or performance loss in human decision-making.
@@
We argue for the identification of a minimum set of theoretical commitments for the task induced by a study of human decision-making from information displays to be “well-defined:” that is, for it to be possible to establish normative behavior. Fortunately, existing theories take on this challenge. Statistical decision theory and expected utility theory (Savage, 1972; Steele and Stefánsson, 2015; Von Neumann and Morgenstern, 2007) provide a rigorously derived and widely applicable framework for studying decision-making. Information economics formalizes information structures for decision problems(Bergemann and Morris, 2019), which may by induced by design choices about how to visualize or explain model predictions.
@@
We synthesize a widely applicable definition of a well-defined decision problem from statistical decision theory and information economics, and motivate the value of this approach to HCAI and related research concerning decision-making from data-driven interfaces. Our first contribution is to establish and motivate a minimum set of components that must be defined for a decision problem to allow identification of an optimal decision so as to identify bias in human decisions. Using the concept of a rational Bayesian agent, we show how performance loss can only be conceived of when a study participant could in theory identify the normative decision from the information they are provided. We find that while 35 of a sample of 46 existing studies draw conclusions about shortcomings in human decisions with prediction displays, only for 6 (17%) of these is the study task well-defined in that participants are given sufficient information to identify, at least in principle, the best decision. We use examples to illustrate the epistemological risks to such conclusions and provide recommendations for experimenters to improve the interpretability of their study results.
@@ 2. Defining a Decision Problem
We define a decision problem and corresponding standard for optimal behavior so as to identify performance loss relative to that standard. Our definition is intended for controlled evaluation of human behavior, a.k.a. normative decision research. Such evaluative studies require that ground truth can be identified for any states that the study participant is queried about. Behavioral data (which may be produced by humans or by simulation) is collected under controlled conditions, with the goal of learning about behavior induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank different assistive elements by human performance (e.g, different visualizations (Correll et al., 2018) or AI explanation strategies (Liu et al., 2021)), or to test a hypothesis about how humans make decisions or what will help them do better (e.g., cognitive forcing functions will improve AI-assisted decisions (Buçinca et al., 2021)).
@@ 2.1. Decision Problem
When we present a person (hereafter agent) with a decision problem, we give them a choice of response (action) $a$ from some set of possible responses (action space) $A$. In decision tasks corresponding to prediction tasks, the action space is a probabilistic belief over the state space, i.e., $A=\Delta(\Theta)$. The agent faced with a decision task must consider the probability of different states induced by the signal, and choose the best action under the scoring rule, a function that assigns a score representing the quality of the action for the realized state.
@@ 2.2. Normative Behavior
where $q$ describes the agent’s belief distribution, the probability distribution over the states that the agent believes the states of the world are drawn from.
@@ 2.3. Example
2.4. Scoring Rules
Critical to a decision problem is a scoring rule $S$ : $A\times\Theta\to{\mathbb{R}}$, a function that maps the action and state to a quality or payoff. A scoring rule has two possible uses: to incentivize participants and to evaluate the quality of the decision responses.
@@ 2.4. Scoring Rules
Whether the researcher uses the same scoring rule for both incentives and evaluation, or two different rules, has important implications for interpreting the results. For the recidivism example, if the researchers analyzed the collected responses using the above rule (i.e., reporting performance in payoff space) in addition to communicating it to participants for incentivization, then the utility associated with performance is defined by a single scoring rule, and the interpretation of a “good” decision is consistent across incentivization and evaluation. If we imagine instead that the researchers decided to consider errors on a more equal footing in analyzing their data, for example reporting the results in terms of accuracy, where each type of error is associated with the same cost, the meaning of a given action differs between incentivization and evaluation. Because a scoring rule induces a meaning on actions by defining a range of beliefs for which an action is optimal, a challenge to interpretation arises whenever two different rules are used. One rule will by definition “misinterpret” the actions of the other by implying a different normative standard.
An exception to the problems induced by using two different rules occurs when the rule used for incentivization has certain properties. Specifically, among scoring rules for beliefs, we can distinguish those that are proper scoring rules: scoring rules for which the optimal action is to predict the true distribution, i.e., $p\in\operatorname*{arg\,max}_{a\in A}S(a,p)$. When a proper scoring rule is used, the researcher can assume that an agent who understands the rule properly will report their true beliefs, as it is not possible to score higher by reporting some alternative beliefs. These beliefs can then be plugged into a different scoring rule without the risk of misinterpretation.
@@
In practice, decision problems with small actions spaces may be less subject to these challenges. Applying the construction in Equation 6 to a binary decision gives a threshold rule where for each state, all reported probabilities below a threshold receive one score and those above receive another score. Since the payoff depends only weakly on the belief, such scoring rules may be less sensitive to the agents’ understanding of the signal. For this reason, including a decision problem that directly also elicits beliefs using a proper-scoring rule is a good idea whenever eliciting responses using an action space that is not over beliefs, as it enables plugging the elicited beliefs into other scoring rules. [^1] See, e.g., Wu et al. (2023)’s analysis of (Kale et al., 2020).
In general, the choice of possible scoring rules will depend on the nature of the action space. If, for example, the researcher designed the recidivism decision study to elicit the agent’s prediction of the probability of recidivism, they could employ the quadratic scoring rule (i.e., squared loss) to score the accuracy of the probabilistic beliefs. In general, a coarser action space (i.e., with fewer choices available to participants) will produce less informative responses for the purpose of resolving observed performance relative to a normative standard.
2.5. Learning and Feedback
Given a definition of normative behavior dictated by Bayesian decision theory, we would expect the decision-maker to update their beliefs given any information that informs on the joint distribution between signals and states. One source of such information is showing the agent the realization of the state after they make their choice. Imagine that participants in the recidivism decision study see feedback (i.e., are told the realized state, whether or not the defendant recidivated) after each decision they make. Assuming the signal is not revealing of the posterior probability of the state (recidivate or not), we would naturally expect the participants to learn about $\pi$ from this feedback as they complete trials, affecting their posterior beliefs (Equations 3 and 4).
2.6. Decision-Making Under Risk and Certainty
The phrase “decision-making under uncertainty” is defined in the economics literature as the case where the decision-maker is not directly given the probability distribution over the state but instead must infer it. Alternatively, decision-making under risk describes the situation when this distribution is assumed to be known to the decision-maker (Steele and Stefánsson, 2015). Under such an assumption we simply replace $q$ in Equation 2 with the known distribution.
@@ 3. Defining Bias in Responses to a Decision Problem
The primary motivations for using the above framework are epistemological in nature: they concern what we can know about the results of an experiment. To interpret the responses to a decision problem in an experiment on human decision-making as evidence of a faulty decision process, the experiment must provide the participant with enough information to in principle identify the normative decision that will be used to judge their behavior. In other words, does the experiment give the participant enough information to align their understanding of the decision problem with its normative interpretation?
@@
When an experiment does not provide sufficient information for the participant to solve the task in principle, there is the potential for multiplicity concerning the data-generating model. Such multiplicity has implications for interpreting the study results. As an example, consider the common approach to defining appropriate reliance on AI as agreeing with the AI when it is correct and disagreeing when it is not (Bansal et al., 2021; Gonzalez et al., 2020; Wang and Yin, 2021; Yang et al., 2020). Imagine a user receives information about the model’s overall performance unconditional on the predicted class. Consider two data generating models that are consistent with the provided information: one where errors are always false positives, one where errors are always false negatives. Imagine a user who can only correct false positives. If we observe the user disagreeing with the model prediction some observed proportion of the time, we would say they are overrelying under the first data generating model, but underrelying under the second. However, both are compatible with the information they were given.
@@
It is worth noting that even if the researchers do not think about the elements of their experiment in terms of the components identified above (i.e., statespace, data-generating model, signaling policy, action space, scoring rule), the design of their study still induces a decision problem whenever two simple conditions hold (Wu et al., 2023). First, there must be some uncertainty about the state (i.e., there is more than one data scenario used to generate stimuli) [^3] A single trial study where all participants view the same signal would not qualify., and second, the researchers believe it is possible to provide a scoring rule that is consistent across participants (eliminating, for example, studies of emotional responses).
@@
When these conditions are met, we can characterize forms of bias as loss in decision performance according to four possible sources, which we next define.
@@ 3.1. Defining Sources of Loss
Prior loss (1). The agent must obtain prior beliefs that match the prior beliefs the researchers assume in analyzing the problem. Prior loss is the loss in performance due to the difference between the agent’s prior beliefs and those used by the researchers to calculate the normative standard. Explicit specification of priors is rare in decision studies used to evaluate interfaces in practice. The consequences of this vary depending on the nature of the signaling strategy. When the signal is fully revealing of the state (i.e., the posterior is a point mass), or directly provides the posterior probability of the state, there is no need for the prior. However, whenever the signal cannot directly provide the probability of the state, the prior must be communicated for a well-defined decision problem.
@@
Optimization loss (4). The agent decides by choosing a good action under their posterior beliefs. Optimization loss is the loss in performance due to not identifying the optimal action under the scoring rule. Optimization loss may include what psycho-physicists call “lapse” responses (e.g., from fatigue or inattention), a catch-all term used to denote responses that do not necessarily reflect the underlying function capturing the person’s beliefs. Or it may include “elicitation errors,” where what participants report on the response scale does not reflect what they had in mind as a result of bias or imprecision induced by the elicitation method.

### v3

@@ Abstract.
How well people use information displays to make decisions is of primary interest in human-centered AI, model explainability, data visualization, and related areas. However, what constitutes a decision problem, and what is required for a study to establish that human decisions could be improved remain open to speculation. We propose a widely applicable definition of a decision problem synthesized from statistical decision theory and information economics as a standard for establishing when human decisions can be improved in HCI. We argue that to attribute loss in human performance to forms of bias, an experiment must provide participants with the information that a rational agent would need to identify the utility-maximizing decision. As a demonstration, we evaluate the extent to which recent evaluations of decision-making from the literature on AI-assisted decisions achieve this criteria. We find that only 10 (26%) of 39 studies that claim to identify biased behavior present participants with sufficient information to characterize their behavior as deviating from good decision-making in at least one treatment condition. We motivate the value of studying well-defined decision problems by describing a characterization of performance losses they allow us to conceive. In contrast, the ambiguities of a poorly communicated decision problem preclude normative interpretation. We conclude with recommendations for practice.
@@ 1. Introduction
Research areas in HCI and related fields that rely heavily on decision-making lack a consensus on what a decision is, and how decisions are a distinct form of human behavior compared to tasks that are not considered decisions. Human decisions from prediction and other information displays are often evaluated in studies with the goal of identifying forms of bias: ways in which study participants’ behavior deviates from expectations of good decision-making. However, reviewing the range of tasks that get filed under decision-making, it is difficult to ascertain what common task structure makes these studies about decision-making, much less what properties a study needs to have in order for it to be capable of distinguishing human decisions as flawed.
@@
As examples of the inconsistency and underdefinition of good decision-making in the HCI literature, colloquial definitions describe a decision task as a choice between alternatives, or a choice where the stakes are high (such as deciding which of multiple hurricane forecasts is more accurate (Padilla et al., 2015), or whether to screen child welfare cases (Zytek et al., 2021)), neither of which isolates a clear set of assumptions that must hold for a judgment to qualify as a decision. Other accounts cast decision-making as a higher-level task akin to sensemaking (Dimara and Stasko, 2021). For some, the point of studying human decisions is to compare them to a definition of idealized behavior. However, sometimes researchers define this standard in a way that is common across studies (e.g., a better forecast comes closer to the true probability of the event (Green and Chen, 2019; Fogliato et al., 2021)) while other times different definitions of the same concept are used (e.g., appropriate reliance as agreeing with an AI when it is correct and disagreeing when it’s not (Bansal et al., 2021; Buçinca et al., 2021) versus providing a higher subjective rating of trust when the model is correct (Bussone et al., 2015) versus choosing the better of the human versus the AI prediction to bet on for a future task (Dietvorst et al., 2015)). Sometimes the point is to study seemingly “subjective decisions,” like how people react to music recommendations (Kulesza et al., 2012), or decide whether a house seems to match its valuation (Prabhudesai et al., 2023), where it can seem unclear whether normative interpretations of task performance can be applied. Not only is a more coherent definition of decision needed, as recent survey papers on decision making in visualization and human-centered AI call for (Dimara and Stasko, 2021; Lai et al., 2021), but such a definition must clearly define requirements for identifying forms of bias or performance loss in human decision-making.
@@
We argue for the identification of a minimum set of theoretical commitments for the task induced by a study of human decision-making from information displays to be well-defined. By well-defined we mean that it is possible to establish normative behavior, and consequently to conclude that observed human decisions are flawed in some way. We draw on statistical decision theory and expected utility theory (Savage, 1972; Steele and Stefánsson, 2015; Von Neumann and Morgenstern, 2007), which provide a rigorously derived and widely applicable framework for studying decision-making. Information economics formalizes information structures for decision problems(Bergemann and Morris, 2019), which may by induced by design choices about how to visualize or explain model predictions.
@@
We contribute a widely applicable definition of a well-defined decision problem, and an argument for the value of this approach to HCAI and related research concerning decision-making from data-driven interfaces. Our first contribution is to establish and motivate a minimum set of components that must be defined for a decision problem to allow identification of an optimal decision (or set of decisions) so as to identify bias in human decisions. Using the concept of a rational Bayesian agent, we show how performance loss can only be conceived of when a study participant could in theory identify how to optimize their behavior (i.e., identify the normative decision) from the information they are provided. We find that while 39 of a sample of 46 existing studies draw conclusions about shortcomings in human decisions with prediction displays, only for 10 (26%) of these is the study task well-defined in at least one treatment condition, in the sense that participants are given sufficient information to identify, at least in principle, the best decision. We use examples to illustrate the epistemological risks to such conclusions and provide recommendations for experimenters to improve the interpretability of their study results.
@@ 2. Defining a Decision Problem
Our definition of a decision problem and corresponding standard for optimal behavior is intended for controlled evaluation of human behavior, a.k.a. normative decision research. Behavioral data (which may be produced by humans or by simulation) is collected under controlled conditions, with the goal of learning about behavior induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank different assistive elements by human performance (e.g, different visualizations (Correll et al., 2018) or AI explanation strategies (Liu et al., 2021)), or to test a hypothesis about how humans make decisions or what will help them do better (e.g., cognitive forcing functions will improve AI-assisted decisions (Buçinca et al., 2021)). Critically, such evaluative studies require that ground truth can be identified for any states that the study participant is queried about.
@@ 2.1. Decision Problem
When we present a person (hereafter agent) with a decision problem, we give them a choice of response (action) $a$ from some set of possible responses (action space) $A$. In decision tasks corresponding to prediction tasks, the action space is a probabilistic belief over the state space, i.e., $A=\Delta(\Theta)$. The agent faced with a decision task must consider the probability of different states induced by the signal, and choose the best action under the scoring rule, a function that assigns a score representing the quality of the action for the realized state: $S:\Theta\times A\rightarrow\mathbb{R}$.
@@ 2.2. Normative Behavior
where $q$ describes the agent’s belief distribution, the probability distribution over the states that the agent believes the states of the world are drawn from, and $S(a,\theta)$ represents the score assigned by $S$.
@@ 2.3. Example
2.4. Additional Remarks on Scoring Rules
As previously mentioned, a critical yet often poorly understood component of a decision problem is a scoring rule $S$ : $A\times\Theta\to{\mathbb{R}}$ that maps the action and state to a quality or payoff. A scoring rule has two possible uses: to incentivize participants and to evaluate the quality of the decision responses.
@@ 2.4. Additional Remarks on Scoring Rules
Whether the researcher uses the same scoring rule for both incentives and evaluation versus two different rules has important implications for interpreting the results. For the recidivism example, if the researchers analyzed the collected responses using the above rule (i.e., reporting performance in payoff space) in addition to communicating it to participants for incentivization, then the utility associated with performance is defined by a single scoring rule, and the interpretation of a “good” decision is consistent across incentivization and evaluation, which is clearly beneficial for interpreting the results.
If we imagine instead that the researchers decided to consider errors on a more equal footing in analyzing their data, for example reporting the results in terms of accuracy, where each type of error is associated with the same cost, the meaning of a given action differs between incentivization and evaluation. A scoring rule induces a meaning on actions by defining a range of beliefs for which an action is optimal. Consequently, interpreting results is difficult whenever two different rules are used. One rule will by definition “misinterpret” the actions of the other by implying a different normative standard.
One way around this problem when using two different rules is to make sure that the rule used for incentivization has certain properties. Specifically, among scoring rules for beliefs, we can distinguish those that are proper scoring rules: scoring rules for which the optimal action is to predict the true distribution, i.e., $p\in\operatorname*{arg\,max}_{a\in A}S(a,p)$. When a proper scoring rule is used, the researcher can assume that an agent who understands the rule properly will report their true beliefs, as it is not possible to score higher by reporting some alternative beliefs. These beliefs can then be plugged into a different scoring rule without the risk of misinterpretation.
@@
Applying the construction in Equation 6 to a binary decision gives a threshold rule where for each state, all reported probabilities below a threshold receive one score and those above receive another score. Since the payoff depends only weakly on the belief, such scoring rules may be less sensitive to the agents’ understanding of the signal. For this reason, including a decision problem that directly also elicits beliefs using a proper-scoring rule is a good idea whenever eliciting responses using an action space that is not over beliefs, as it enables plugging the elicited beliefs into other scoring rules. [^1] See, e.g., Wu et al. (2023)’s analysis of (Kale et al., 2020).
The choice of possible scoring rules depends on the nature of the action space. In general, a coarser action space (i.e., with fewer choices available to participants) will produce less informative responses for the purpose of resolving observed performance relative to a normative standard. If, for example, the researcher designed the recidivism decision study to also elicit the agent’s prediction of the probability of recidivism, each trial would provide more information, and they could employ a proper scoring rule such as squared loss (i.e., Brier or quadratic rule) or log score to score the accuracy of the probabilistic beliefs. Different proper scoring rules imply different expectations about the set of possible decisions (Levinstein, 2017).
2.5. Extensions
Learning and Feedback Given a definition of normative behavior dictated by Bayesian decision theory, we would expect the decision-maker to update their beliefs given any information that informs on the joint distribution between signals and states. One source of such information is showing the agent the realization of the state after they make their choice. Imagine that participants in the recidivism decision study see feedback (i.e., are told the realized state, whether or not the defendant recidivated) after each decision they make. Assuming the signal is not revealing of the posterior probability of the state (recidivate or not), we would naturally expect the participants to learn about $\pi$ from this feedback as they complete trials, affecting their posterior beliefs (Equations 3 and 4).
Decision-Making Under Risk and Certainty The phrase “decision-making under uncertainty” is defined in the economics literature as the case where the decision-maker is not directly given the probability distribution over the state but instead must infer it. Alternatively, decision-making under risk describes the situation when this distribution is assumed to be known to the decision-maker (Steele and Stefánsson, 2015). Under such an assumption we simply replace $q$ in Equation 2 with the known distribution.
@@ 3. Defining Bias in Responses to a Decision Problem
The primary motivations for using the above framework are epistemological in nature: they concern what knowledge can be gained from the results of a study. Our primary claim in this work is that to interpret the results of an experiment on human decision-making as evidence that human decisions are flawed in some way, the experiment must provide the participant with enough information to in principle identify the normative decision that will be used to judge their behavior. In other words, does the experiment give the participant enough information to align their understanding of the decision problem with its normative interpretation?
@@
When an experiment does not provide sufficient information for the participant to solve the task in principle, there is the potential for multiplicity concerning the data-generating model. Such multiplicity has implications for interpreting the study results. As an example, consider the common approach to defining appropriate reliance on AI as agreeing with the AI when it is correct and disagreeing when it is not (Bansal et al., 2021; Gonzalez et al., 2020; Wang and Yin, 2021; Yang et al., 2020). Imagine a user receives information about the model’s overall performance unconditional on the predicted class. Consider two data-generating models that are consistent with the provided information: one where errors are always false positives, one where errors are always false negatives. Imagine a user who can only correct false positives. If we observe the user disagreeing with the model prediction some observed proportion of the time, and getting some of those instances wrong we would say they are overrelying under the first data-generating model. If they disagree with some model predictions under the second data-generating model, we would say they are underrelying. However, both are compatible with the information they were given.
3.1. Scope
@@
It is worth making several remarks on the scope of our argument. It is important to recognize that even if the researchers do not think about the elements of their experiment in terms of the components identified above (i.e., statespace, data-generating model, signaling policy, action space, scoring rule), the design of their study still induces a decision problem whenever two simple conditions hold (Wu et al., 2023). First, there must be some uncertainty about the state (i.e., there is more than one data scenario used to generate stimuli) [^3] A single trial study where all participants view the same signal would not qualify., and second, the researchers believe it is possible to provide a scoring rule that is consistent across participants (eliminating, for example, studies of emotional responses).
It should also be noted that our argument is that the interpretation of results is restricted when a well-defined decision problem is not conveyed to participants. Researchers can certainly conduct and report on studies that do not conform to this approach, but they cannot conclude that people made bad decisions if they were not given sufficient information to know what a good decision is. For example, imagine a study comparing the decision performance of users with and without the use of an AI, where no prior is given to participants, nor a scoring rule. Imagine that the performance measured in terms of accuracy is higher access to an AI. Given an appropriately designed randomized experiment the researcher could conclude that there is a causal effect of the AI on decisions. However, they could not claim that people are making worse or biased decisions without the AI, because they would be scoring their decisions using information they never provided to participants. When we say that one condition did better, we should acknowledge that we did not give them a chance to fully understand and optimize their behavior for the problem that we analyzed.
Finally, a common misinterpretation of the use of a rational agent model is that we are implying that human behavior is expected to be perfectly rational, or approximately rational. This is not the case. While a rational agent approach formally justifies what it means to best use information to inform the selection of an action using statistical decision theory, participants in our experiment may use different strategies. The point is that without a well-defined normative response dictating what it means to best respond to the information we have provided, we are hard-pressed to evaluate participant behavior. Even if we do not expect this procedure to characterize what human agents do, it provides a basis for interpreting our results via deviations from a well-understood standard.
@@
We we describe below, when the conditions above are met, we can characterize forms of bias as loss in decision performance according to four possible sources.
@@ 3.2. Defining Sources of Loss
Prior loss (1). The agent must obtain prior beliefs that match the prior beliefs the researchers assume in analyzing the problem. Prior loss is the loss in performance due to the difference between the agent’s prior beliefs and those used by the researchers to calculate the normative standard. Explicit specification of priors is rare in decision studies used to evaluate interfaces in practice. The consequences of this vary depending on the nature of the signaling strategy. When the signal is fully revealing of the state (i.e., the posterior is a point mass indicating the state with certainty), or directly provides the posterior probability of the state, there is no need for the prior. However, whenever the signal cannot directly provide the probability of the state, the prior must be communicated for a well-defined decision problem.
@@
Optimization loss (4). The agent decides by choosing a good action under their posterior beliefs. Optimization loss is the loss in performance due to not identifying the optimal action under the scoring rule. Optimization loss may include what perceptual psychologists call “lapse” responses (e.g., from fatigue or inattention), a catch-all term used to denote responses that do not necessarily reflect the underlying function capturing the person’s beliefs. Or it may include “elicitation errors,” where what participants report on the response scale does not reflect what they had in mind as a result of bias or imprecision induced by the elicitation method.